Are the following statements true?
32 m : 64 m = 6 sec : 12 sec 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

32 m : 64 m = 6 sec : 12 sec
32 / 64 = 1 / 2
6 / 12 = 1 / 2
Hence, True

Thanks